Let V= R^3

W={(x1, x2, x3)| x1-x2 =x3}. Show that W is a

subspace of V. Further,find a basis for W and hence,find the dimension of W.


1
Expert's answer
2021-02-25T05:05:56-0500

Let V= R^3

"W={(x_1, x_2, x_3)| x_1-x_2 =x_3}" . Show that W is a

subspace of V.

Further,find a basis for W and hence,find the dimension of W.


To show that W is a subspace of V we need to show the following:

let "r= (r_1, r_2, r_3)" , "q=(q_1, q_2,q_3) \isin W" and "\alpha\isin" field

"\implies q_1-q_2=q_3" and "r_1- r_2= r_3"

  1. W is not empty
  2. for r,q "\isin" W , r+q "\in" W
  3. "\alpha r\isin W"

if "(0, 0,0)\isin" W then "0-0=0"

therefore, W is not empty


"r+q=(r_1, r_2, r_3) +(q_1, q_2,q_3)={ (r_1+q_1), (r_2+q_2),(r_3+q_3)}"

since "(r_1+q_1)-(r_2+q_2)=(r_1-r_2) +(q_1-q_2)= r_3+ q_3"

"\implies r+q\isin W"


"\alpha r=\alpha (r_1, r_2, r_3)=r= (\alpha r_1, \alpha r_2,\alpha r_3)"

then "\alpha r_1-\alpha r_2=\alpha(r_1-r_2)=\alpha r_3"

"\implies \alpha r\isin W"

therefore W is a subspace of V


"r=\begin{pmatrix}
 r_1 \\
 r_2 \\r_3 
\end{pmatrix} =\begin{pmatrix}
 r_1 \\
 r_2 \\r_1-r_2 
\end{pmatrix} = r_1\begin{pmatrix}
 1 \\
 0 \\1
\end{pmatrix}+r_2\begin{pmatrix}
 0 \\
 0 \\-1
\end{pmatrix}"


"\therefore ,

\begin{bmatrix}
 \begin{pmatrix}
 1 \\
 0 \\1
\end{pmatrix} & \begin{pmatrix}
 0 \\
 0 \\-1 \end{pmatrix}
 
\end{bmatrix}" is a Basis of W


The Dimension of W is 2